OpenStudy (anonymous):

Solve: (4x^2 - 9)^2 - 11(4x^2 - 9) + 24 = 0.

OpenStudy (anonymous):

\[x = \sqrt{3}\] or \[x=-\sqrt{3}\] or \[x=- \sqrt{17}/2\] or \[x= \sqrt{17} /2\]

OpenStudy (anonymous):

@tropical u use da quadratic equation ok i just work it out normally and get 16x^4-116x^2+204

OpenStudy (anonymous):

we had to solve it...

OpenStudy (anonymous):

@ok thanx 4 the info

OpenStudy (amistre64):

u^2 -11u +24 = 0 ; u=(4x^2 -9) (u-3)(u-8)=0 u=3 or u=8 4x^2 -9=3 4x^2 = 12 2x= +-sqrt(12) x= +-2sqrt(3)/2 = +-sqrt(3) 4x^2 -9=8 4x^2 =17 2x=+-sqrt(17) x= +-sqrt(17)/2 Just another way of doing it :)